The hypotenuse of a right triangle is 20 centimeters. The sum of the other two sides is 28 centimeters. Find the length of two sides.
Answer by edjones(8007)

a+b=28
let b=28-a
a^2+(28-a)^2=20^2
a^2+a^2-56a+784=400
2a^2-56a+384=0
2(a^2-28+192)=0
2(a-12)(a-16)=0
a=12, a=16
So the 2 sides are 12 cm and 16 cm
Ed